OpenStudy (anonymous):

a=-3x+7 It says the answer is 7-a/3? Help?

OpenStudy (anonymous):

you are solving for x a-7=-3x make the sign positive for x by multiplying -1 or changing every other sign -a+7=3x x=(7-a)/3